Circular No. 5921

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993aj IN ANONYMOUS GALAXY M. Turatto, Padova; and C. Gouiffes, Saclay, write: “We have obtained

a spectrogram (600-980 nm) of SN 1993aj (cf. IAUC 5915) on Jan. 11.3 UT with the European Southern Observatory 3.6-m telescope (+ EFOSC1) at La Silla. The preliminary reduction shows broad features, confirming that the object, projected on the disk of an edge-on galaxy, is indeed a supernova. Strong emissions are centered at 632.9 and 692.8 nm, and two absorptions are measured at 613.0 and 663.7 nm. The spectrum resembles those of type-Ia supernovae (e.g., SN 1989B, Barbon et al. 1990, A.Ap. 237, 79) at ∼ 2-3 weeks past maximum, once a correction has been made for z ∼ 0.075 — rather than the spectra of type-II supernovae, as suggested on IAUC 5917. Assuming this redshift, the magnitudes reported on IAUC 5919 are consistent with those of a type-Ia supernova in the proximity of maximum. The parent galaxy is a highly inclined early-type spiral. In the observed range, its faint spectrum shows no strong features. A weak absorption at 634.6 nm corresponds to the Na I D line if z = 0.075 is adopted.”

GRO J0422+32 M. R. Zapatero-Osorio, R. Rebolo, M. Kidger, M. Serra-Ricart, and

E. L. Martin, Instituto de Astrofisica de Canarias (IAC), report: “A new mini-outburst of GRO J0422+32 has been detected using the 0.80-m IAC telescope at the Teide Observatory. CCD observations yield the following magnitudes: Jan. 5.99 UT, R = 19.1; 7.13, R = 17.9; 7.84, R = 17.5; 7.96, B = 18.0, V = 17.6, R = 17.4, I = 16.6; 8.10, B = 17.9, V = 17.4, R = 17.2, I = 16.4; 8.97, R = 17.1; 11.04, R = 17.1.”

1993 VW R. P. Binzel and S. J. Bus, Massachusetts Institute of Technology,

report spectroscopic observations of 1993 VW obtained on Jan. 7 UT with the 2.4-m Hiltner telescope of the Michigan-Dartmouth-MIT Observatory. Measurements over 0.4-1.0 μm show spectral properties resembling the Qtype asteroid (1862) Apollo and possibly ordinary chondrite meteorites. Additional physical measurements are recommended. Note that a favorable apparition occurs in May 1994.

1994 January 13 Daniel W. E. Green